OpenStudy (anonymous):

Mia's average driving speed is 10 mph than thelmas. In the same length of time it takes mia to drive 240 miles, thelma only drives 200 miles. What is mias average speed

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

I need the answer right now! I just can't figure it out!

OpenStudy (anonymous):

x(240/(x+10))=200 240x = 200x + 2000 40x = 2000 x= 50 so Mia's driving speed is 60mph

OpenStudy (anonymous):

let X = Thelma's speed. X+10 is Mia's speed then. T is the amount of time that passes. (X+10)T=240 for Mia XT=200 for Thelma Solve for T for either Mia or Thelma and then substitute that into the other equation.

OpenStudy (anonymous):

Refer to the Mathematica 9 attachment.